Use technology to find the​ P-value for the hypothesis test described below. The claim is that for a smartphone​ carrier's data speeds at​ airports, the mean is mu equals11.00 Mbps. The sample size is nequals29 and the test statistic is t equals negative 1.989.

​P-value =  ​(Round to three decimal places as​ needed.)
Answer #1

Solution :-

Given :-

t statistic = - 1.989

Sample Size ( n ) = 29

df = n - 1

df = 29 - 1

df = 28

Claim - for a smartphone carrier data speeds at airport the mean u = 11 Mbps

Hypothesis :-

Ho : u = 11

H1 : u 11

The Given test is two tailed test.

P -Value is,

By using t value table,

at , t= - 1.989 and df = 28 ( for two tailed test )

P -Value = 0.057